### v2

# Can Few-shot Work in Long-Context? Recycling the Context to Generate Demonstrations

Abstract
Despite recent advancements in Large Language Models (LLMs), their performance on tasks involving long contexts remains sub-optimal. In-Context Learning (ICL) with few-shot examples may be an appealing solution to enhance LLM performance in this scenario; However, naïvely adding ICL examples with long context introduces challenges, including substantial token overhead added for each few-shot example and context mismatch between the demonstrations and the target query. In this work, we propose to automatically generate few-shot examples for long context QA tasks by recycling contexts. Specifically, given a long input context (1-3k tokens) and a query, we generate additional query-output pairs from the given context as few-shot examples, while introducing the context only once. This ensures that the demonstrations are leveraging the same context as the target query while only adding a small number of tokens to the prompt. We further enhance each demonstration by instructing the model to explicitly identify the relevant paragraphs before the answer, which improves performance while providing fine-grained attribution to the answer source. We apply our method on multiple LLMs and obtain substantial improvements (+23% on average across models) on various QA datasets with long context, especially when the answer lies within the middle of the context. Surprisingly, despite introducing only single-hop ICL examples, LLMs also successfully generalize to multi-hop long-context QA using our approach.
[…]
tcb@breakable
Can Few-shot Work in Long-Context?
Recycling the Context to Generate Demonstrations
Arie Cattan1,2thanks: Work done during an internship at Google.  Alon Jacovi2  Alex Fabrikant3  Jonathan Herzig2 Roee Aharoni2Hannah Rashkin3Dror Marcus2 Avinatan Hassidim2Yossi Matias2Idan Szpektor2Avi Caciularu2 1Bar-Ilan University  2Google Research  3Google DeepMind cattana@google.com
1 Introduction
Figure 1: Performance of Gemini Pro (v1.0) on a sample of the Lost-in-the-middle dataset (Liu et al., 2023). The X-axis is the position of the relevant passage in the context. The baseline (blue line) displays a U-shaped curve, performing well only when the relevant passage is at the beginning or end of the input. The oracle (green line) shows significant performance gain when the relevant passage ID is provided in the prompt, showing that the main bottleneck is the identification of supporting evidence(s). DoubleDipper (our method, orange line) flattens this U-shaped trend.
Long contexts are prevalent in various domains, ranging from legal documents and scientific articles to lengthy reports and novels. These may consist of a single extensive document or multiple passages, typically retrieved through specific retrieval mechanisms (e.g., RAG; Lewis et al., 2020).
Yet, while Large Language Models (LLMs) have demonstrated impressive capabilities in a variety of tasks including answering questions requiring one or multiple reasoning steps, they often struggle to answer simple questions when faced with long contexts. Despite substantial engineering efforts (Chen et al., 2023) to extend the context window of LLMs to extremely long inputs (32k and even 1M tokens), these models continue to struggle with much shorter inputs, comprising only a few thousand tokens.
In order to answer questions from long inputs, models should implicitly identify relevant information segments and then reason over these segments to formulate an answer. It has been shown that LLMs struggle when the relevant information is buried in the middle of the context (Liu et al., 2023) or obscured by numerous irrelevant details (Levy et al., 2024). Our analysis (illustrated in Figure 1) identifies the identification of relevant information as a primary performance bottleneck of current models in long contexts.
In this work, we introduce a novel method to enhance the QA performance of LLMs in long input setups (to allow direct comparisons across a wide swath of models, we limit "long context" here to 1-3k tokens). Our approach, termed DoubleDipper, leverages LLMs’ In-Context Learning (ICL) capability and is based on two principles. First, instead of typical ICL, where each few-shot example is standalone with a separate length context and a question-answer (QA) pair, we propose to recycle the given input context and automatically generate few-shot examples from this context. Specifically, we randomly select a few paragraphs from the given input context and generate QA pairs for each passage. These generated QAs serve as demonstration examples and are placed between the input context and the target input question. Figure 2 illustrates the differences between the traditional ICL with few-shot examples and DoubleDipper. Second, we further enrich each ICL example to instruct the model to explicitly identify the paragraph(s) containing the relevant information before generating the answer. This can be regarded as a structured Chain of Thought that incentivizes the model to pinpoint relevant information before reasoning, an essential capability for long-context processing.
Figure 2: Comparison of traditional In-Context-Learning (ICL) and our new method. In traditional ICL (left), each example comprises a possibly lengthy context, accompanied by a query and an answer, typically derived from the training dataset. Conversely, our approach (right) simplifies each example to just a question and an answer, both of which are generated directly from the provided input context.
By generating few-shot demonstrations from various sections of the input context while instructing the model to identify relevant passages, DoubleDipper encourages the model to develop deeper reading comprehension skills specific to the given input evidence. This, in turn, allows the model to answer subsequent queries with higher accuracy. DoubleDipper presents several advantages. In terms of efficiency, since each example does not include its own input context, our method adds to the original prompt a minimal number of tokens, resulting in a substantially cheaper inference than traditional ICL. Additionally, recycling the same context for ICL demonstrations ensures that the few-shot examples refer to the same domain as the input question, thus obviating the need for external retrieval processes. Finally, DoubleDipper generates answers with attribution to relevant paragraphs, improving the model’s lookup ability and offering transparency, which substantially simplifies human evaluation (Slobodkin et al., 2024).
We applied DoubleDipper to a variety of LLMs, both commercial (Gemini Pro and Gemini Ultra; Team et al., 2023) and open-source (Llama (Touvron et al., 2023), Mistral (Jiang et al., 2023) and Gemma (Team et al., 2024)), and evaluate it on various QA datasets with long inputs, including common multi-hop QA datasets. Our experiments demonstrate that with only 3 self-generated few-shot examples, DoubleDipper consistently outperforms the baseline on our evaluation set by 23% on average across models. In addition, for some models, DoubleDipper enhances the robustness to the position of the relevant information within the text. Interestingly, while our few-shot examples focus on single-paragraph answers, DoubleDipper generalizes well to multi-hop QAs and where the answer requires information from multiple passages.
2 Background
Challenges in Long Context for Language Modeling.
LLMs have been well-documented to struggle when input length grows (An et al., 2023), and especially so when it exceeds input lengths seen during training (Anil et al., 2022). Various methods have been proposed to advance long-context capabilities: Architectural, e.g., to augment the embedding layer to cleverly extrapolate to unseen lengths (Vaswani et al., 2017; Press et al., 2021; Caciularu et al., 2022); and via data, e.g., to incorporate longer inputs and more challenging long-context scenarios into training (He et al., 2023; Chen et al., 2023). However, this challenging problem stubbornly remains in competitive models today (Liu et al., 2023; Bishop et al., 2023; Levy et al., 2024). In contrast to the above methods, DoubleDipper does not involve training or architectural changes.
In documenting and exploring LLM performance in long-context settings, many different benchmarks targeting it have been proposed, such as Scrolls and Zero-Scrolls (Shaham et al., 2022, 2023), Loogle (Li et al., 2023), LongBench (Bai et al., 2023), inter alia. We describe the most relevant evaluation benchmarks used in this work in Section 4.
In-Context Learning
The area of in-context learning (ICL) is a class of prompting techniques where demonstrations are added to the prompt in order to steer or improve model behavior (Min et al., 2022a). Typically, in-context learning involves hand-crafted demonstrations (Song et al., 2022), automatic retrieval of demonstrations from a larger set (Paranjape et al., 2023), or instructing the model to perform various tasks one after another in a pipeline (Gao et al., 2022). Recent improvements in long-context capabilities of LLMs have also had effect on improving the yield from in-context learning by simply using more short-length demonstrations (Agarwal et al., 2024).
While such methods are widely used for their effectiveness (Brown et al., 2020b; an Luo et al., 2024), they remain under-explored in settings of long-context. The reason is simple: If the context is already extremely long, adding additional demonstrations comparable in length to the input context will likely amplify the existing limitations of long context handling (Li et al., 2024). In this work, we tackle these challenges and present a novel ICL method for long contexts.
3 DoubleDipper
Figure 3: Example of DoubleDipper applied to the MuSique dataset. Given 20 passages as input, DoubleDipper randomly selects 3 passages (specifically passages 15, 5, 17) and automatically generates a question-pair for each one. As each QA is associated with its respective paragraph, we form the demonstrations to instruct the model to identify the relevant passage(s) and the correct answer.
Recall that our work focuses on the task of question answering (QA) with long input context comprising multiple paragraphs. In addition to the answer, we aim to identify the supporting paragraphs in order to provide attribution. Formally, given a long input text $D$ composed of $n$ paragraphs $D=\{p_{1},p_{2},...,p_{n}\}$ and a question $q$, the goal is to generate the answer $a$ and identify the set(s) of paragraphs that support the answer $S=\{s_{1},...,s_{k}\}$. The number of the supporting paragraphs is not known in advance and can be one or more.
We describe DoubleDipper, an efficient method for improving the performance of large language models (LLMs) when dealing with long contexts. The core principles of DoubleDipper involve: (1) recycling the input context to automatically generate few-shot examples, and (2) “teaching” the model via in-context learning (ICL) to explicitly pinpoint the supporting paragraphs before generating the answer.
Figure 3 illustrates DoubleDipper. Starting with the input paragraphs $D$, we initially select $k$ paragraphs at random (e.g., paragraphs $15$, $5$, and $17$, for $k:=3$). For each chosen paragraph, we prompt the model to formulate a question that pertains to the specific paragraph, accompanied by an appropriate answer (for further details on prompt specifications, refer to Appendix A). Each generated QA pair is directly associated with its origin paragraph, enabling us to assemble the following structured in-context demonstration, shown as the DoubleDipper block in Figure 3:
Question $:q_{i}$
Evidence $:p_{i}$
Answer $:a_{i}$
Here, $p_{i}$ indicates the index of the paragraph associated with the QA pair $(q_{i},a_{i})$. Given a test question $q$, we then form a QA prompt by concatenating the original input context $D$, the compiled demonstrations and $q$. The model first generates the one or more indices of the supporting paragraph(s), followed by the answer.
Unlike traditional few shot examples that instruct the model about a specific task, DoubleDipper aims to coach the model on how to “handle” the specific input context. This is achieved by guiding the model to explicitly localize relevant information before generating the answer. By randomly sampling multiple paragraphs from the input, DoubleDipper guarantees that the ICL demonstrations involve reading different parts of the context, allowing the model to better comprehend the input text. Beyond improving the performance of the QA task, instructing the model to provide the supporting paragraphs is valuable on its own as it offers transparency and substantially eases human evaluation (Slobodkin et al., 2024).
DoubleDipper offers several advantages. First, as each example in the demonstration consists only of a question, an answer and the ID of relevant passage, the number of added tokens due to the extra demonstrations is minimal, leading to a low additional cost and computation. Furthermore, by reusing the same context to generate demonstrations, our approach guarantees that all few shot examples are derived from the exact same domain as the input query (Rubin et al., 2022).
Finally, we observe that, although the QA pairs in the demonstration are confined to individual paragraphs, the actual query $q$ may require reasoning over multiple paragraphs (i.e., multi-hop QA). Surprisingly, LLMs can generalize from DoubleDipper local examples to these complex, global questions and successfully generate indices to multiple paragraphs (see Section 5).
4 Experiments
Datasets
Dataset # Instances Avg. # tokens
Lost-in-the-middle 2,500 2,815
FLenQA 1,500 3,225
HotpotQA 500 1,646
2Wiki 500 1,222
MuSiQue 500 2,549
Table 1: Evaluation datasets in our experiments. The average number of tokens is computed according to Gemma’s tokenization of the simple prompt.
We apply our method to various datasets, each presenting its own domain-specific challenges. We selected these datasets because the supporting paragraphs are also annotated. Overall our evaluation set include 5.5K instances, with statistics of each dataset given in Table 1.
The Lost-in-the-middle dataset (Liu et al., 2023) includes examples from NaturalQuestions-Open (Kwiatkowski et al., 2019; Lee et al., 2019). Each instance consists of twenty Wikipedia passages, with only one passage containing the answer to the query. The remaining passages are distractors that are lexically similar but do not contain the answer. To assess the robustness of large language models (LLMs) to the position of relevant information, Liu et al. (2023) evaluated cases where the relevant passage appeared in positions 1, 5, 10, 15, and 20. Following their methodology, we sampled 500 instances for each position, resulting in a total of 2,500 instances.
FLenQA (Levy et al., 2024) is a benchmark that includes simple questions with answers of either “True” or “False” based on two key sentences. FLenQA includes three subtasks. The first subtask is Monotone Relations (MonoRel), where each instance asks whether a transitive relation between two entities holds based on the context (e.g., "Is X younger than Y?" based on the sentences "X is younger than Z" and "Z is younger than Y"). The second subtask, People In Rooms (PIR), involves one key sentence indicating that a person is in a specific room and another key sentence describing a property of this room. The question asks whether the person is in a room with the described property. The final subtask is Simplified Rule Taker (SRT), based on RuleTaker (Clark et al., 2020). Each instance consists of a logical rule, two sentences each introducing a fact, and a question over the rule and facts. For each subtask, FLenQA includes contexts with varying lengths, from 50 to 3,000 tokens, by simply adding irrelevant text, demonstrating consistent performance degradation with increased input length. In our experiments, we sampled 250 instances for each subtask with input lengths of 2,000 and 3,000 tokens, leading to a total of 1,500 instances for FLenQA.
In addition, we evaluate our method on common multi-hop QA benchmarks. We sampled 500 instances from each of the following datasets: HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020), and MuSiQue (Trivedi et al., 2021). In all these datasets, the input text includes multiple passages, and models need to perform at least two steps of reasoning over different passages in order to answer the question.
Models
We apply our method to a variety of models, both commercial and open-source. The commercial models include Gemini-1.0-Ultra and Gemini-1.0-Pro (Team et al., 2023). The open-source models we tested are Llama-2-7b-chat, Llama-2-13b-chat, Llama-2-70b-chat (Touvron et al., 2023), Gemma-2b-it (v1.1), Gemma-7b-it (v1.1)(Team et al., 2024) and Mistral-7b-instruct (v0.2)(Jiang et al., 2023). Details about models’ size and context window are shown in Table 2.
For few-shot generation, we conducted two distinct experiments in order to analyze the effect of different question generation models. In the first experiment, dubbed DoubleDipper (Self), we employed the same LLM for both generating the demonstrations and answering to the query. In the second experiment, named DoubleDipper (PaLM 2), we generated a single set of few-shot examples using PaLM 2 (Anil et al., 2023) and then supplied these examples to various LLMs. This approach allowed us to assess the performance consistency across different models when provided with the same set of demonstrations.
Baselines
We compare DoubleDipper to the vanilla baseline, which takes as input the entire context $D$ and the query $q$ and generate only the answer $a$.
Size Context Window
Gemini v1.0 Pro 32k
Ultra 32k
Llama-2 7B 4k
13B 4k
70B 4k
Mistral 7B 8k
Gemma 2B 8k
7B 8k
Table 2: Models used with DoubleDipper.
Evaluation
We evaluate each dataset with the original evaluation metrics. Namely, we report Accuracy for Lost-in-the-middle (Liu et al., 2023) and FLenQA (Levy et al., 2024), and Token F1 for HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020) and MuSique (Trivedi et al., 2021).
In addition to the task’s accuracy, we also evaluate the performance of the identification of the supporting paragraph(s), by computing the F1 score on the predicted set of supporting passages compared to the ground truth (Yang et al., 2018; Ho et al., 2020; Trivedi et al., 2021).
Implementation Details
We randomly select three passages from the input (see Section 6 for an analysis of the number of self-generated demonstrations on the performance), each containing at least two sentences, and ask the model to generate five QA pairs per passage (see Appendix A for the exact prompt). We then randomly select a single QA pair for each passage to form the few shot demonstrations. For all experiments, including few-shot generation and question-answering, we use a temperature setting of 0.
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro 39.4 49.0 73.0 30.8 54.4 28.1 18.6 22.1
   + DoubleDipper (Self) 50.0 58.4 85.6 28.6 52.2 56.6 36.5 32.0
   + DoubleDipper (PaLM 2) 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemini Ultra 53.9 53.0 81.2 30.8 66.0 61.6 42.6 42.3
   + DoubleDipper (Self) 53.1 57.2 73.4 20.6 54.6 68.9 51.5 45.6
   + DoubleDipper (PaLM 2) 54.7 56.8 78.6 25.0 56.6 69.1 52.6 44.2
Gemma-2b-it (v1.1) 35.9 24.0 53.4 49.8 49.6 34.6 28.1 11.6
   + DoubleDipper (Self) 35.5 25.5 66.0 54.2 49.2 24.7 18.6 10.5
   + DoubleDipper (PaLM 2) 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Gemma-7b-it (v1.1) 27.2 7.1 52.2 49.8 48.6 12.3 16.7 3.4
   + DoubleDipper (Self) 47.6 45.8 92.6 74.2 45.4 40.4 22.4 12.4
   + DoubleDipper (PaLM 2) 51.3 44.0 94.8 74.8 45.2 49.6 32.1 18.4
Mistral-7b-instruct (v0.2) 45.8 61.3 70.6 68.0 49.0 36.3 21.7 13.9
   + DoubleDipper (Self) 48.9 60.4 96.2 84.6 42.6 28.3 17.8 12.6
   + DoubleDipper (PaLM 2) 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Llama-2-7b-chat 31.6 42.0 46.8 39.0 37.4 32.7 15.1 8.2
   + DoubleDipper (Self) 33.3 45.6 64.2 37.2 35.4 23.3 17.8 9.2
   + DoubleDipper (PaLM 2) 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Llama-2-13b-chat 32.4 51.2 54.2 40.6 43.8 18.2 10.8 7.9
   + DoubleDipper (Self) 34.9 46.9 66.0 44.6 38.4 22.7 19.3 8.4
   + DoubleDipper (PaLM 2) 38.3 47.6 71.0 37.8 36.4 32.8 32.2 10.4
Llama-2-70b-chat 45.3 61.1 66.6 76.4 50.2 33.8 17.7 11.0
   + DoubleDipper (Self) 46.1 57.2 85.6 76.6 44.8 27.8 20.3 10.7
   + DoubleDipper (PaLM 2) 51.6 54.2 88.2 78.0 47.0 44.5 34.2 15.0
Table 3: Results of the baseline, DoubleDipper (Self) and DoubleDipper (PaLM 2) on various QA datasets.
5 Results
Table 3 presents the QA performance of our two variants of DoubleDipper, namely DoubleDipper (Self) and DoubleDipper (PaLM 2) on our evaluation set. We report the performance of the supporting paragraphs prediction in Appendix B. On average across models and datasets, our straightforward method, which automatically generates few-shot examples from the provided input context, yields an improvement of 12% for DoubleDipper (Self) and 23% for DoubleDipper (PaLM 2), often by a significant margin (e.g., +15.2 points for Gemini Pro, +7 points for Mistral Instruct, +6.3 points for Llama2 70B). Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique).
Furthermore, following (Liu et al., 2023), Figure 4 shows the performance of the tested LLMs for both the baseline and DoubleDipper (Self) on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer. For Gemini Pro, Gemini Ultra, Gemma 2B, Gemma 7B, Llama 7B, the performance curve for DoubleDipper consistently surpasses the baseline across nearly all document positions. This reveals that beyond improving performance, DoubleDipper can make the model more robust to the position of the relevant document. For Mistral, Llama 2 13B and Llama 2 70B, Figure 4 shows that DoubleDipper performs on par or better than the baseline when the relevant document is in the middle of the input, but still underperforms the baseline at the beginning and the end. This variation can likely be attributed to the inherent biases of LLMs towards the beginning and end of inputs, while adding in context demonstrations mitigates this bias.
Figure 4: Performance (accuracy) of the models on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
6 Analysis
Few-shot vs. Zero-shot
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro Zero-shot 50.0 59.7 77.4 38.8 54.2 51.2 34.6 34.1
DoubleDipper 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemini Ultra Zero-shot 56.5 56.2 76.8 34.8 62.2 68.3 52.4 44.9
DoubleDipper 54.7 56.8 78.6 25.0 56.6 69.1 52.6 44.2
Gemma-2b-it (v1.1) Zero-shot 3.6 6.4 4.2 5.4 0.0 3.3 4.9 0.7
DoubleDipper 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Gemma-7b-it (v1.1) Zero-shot 7.5 2.6 36.0 5.4 0.0 3.9 3.3 1.0
DoubleDipper 51.3 44.0 94.8 74.8 45.2 49.6 32.1 18.4
Mistral-7b-instruct (v0.2) Zero-shot 41.9 59.0 77.4 67.4 51.8 13.7 12.3 11.8
DoubleDipper 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Llama-2-7b-chat Zero-shot 22.4 29.2 45.0 29.2 33.4 9.8 7.9 2.4
DoubleDipper 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Llama-2-13b-chat Zero-shot 14.0 8.0 43.0 3.0 37.0 5.2 0.3 1.2
DoubleDipper 38.0 36.7 71.4 36.6 50.0 27.4 27.4 16.8
Table 4: Comparison of DoubleDipper to prompting models in a zero-shot setting to provide explicitly the supporting paragraphs, then generate the answer.
In order to assess the effectiveness of the few shots in DoubleDipper, we prompt the model in a zero-shot setting to explicitly specify the ID of the relevant passage(s) before generating the answer. The results of the QA task are detailed in Table 4 and the performance in identifying supporting paragraphs is reported in Appendix B.
For all open-source models, this zero-shot prompting strategy leads to a significant decline in QA performance compared to the baseline (see Table 3), which merely prompts the model for the answer (e.g., -32.3 points for Gemma 2B, -9.2 points for Llama 7B, -4.2 points for Mistral). This drop is likely because retrieving relevant passages makes the task more complex compared to standard QA, as models are not typically trained to perform such “retrieval” tasks. This is further supported by the poor performance in supporting passages identification shown in Appendix B, where F1 scores are often close to zero.
Despite these challenges, DoubleDipper’s self-generated demonstrations systematically improve performance for all open-source models on both QA and supporting passages identification across all datasets. On average, Mistral’s QA performance improves by 10.9 points, and its identification of relevant paragraphs increased by 30.5 F1 score. This aligns with previous research suggesting that models can learn new tasks through in-context learning (ICL).
Comparison to traditional ICL
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro ICL 48.6 59.1 69.8 30.0 52.8 55.6 41.0 32.0
DoubleDipper 54.6 58.4 87.6 40.6 52.8 62.2 42.8 38.0
Gemma-2b-it (v1.1) ICL 21.6 0.2 28.2 26.4 13.0 39.7 39.1 4.3
DoubleDipper 41.2 23.6 75.8 57.8 49.4 38.2 28.1 15.4
Llama-2-7b-chat ICL 16.7 0 13.2 13.2 14.2 37.0 38.2 0.7
DoubleDipper 38.5 43.9 67.6 36.8 35.2 39.2 30.1 16.4
Mistral-7b-instruct ICL 52.1 68.5 79 67.8 39.2 49.4 36.7 24.1
DoubleDipper 52.8 58.9 95.6 82.8 44.4 43.1 26.4 18.4
Table 5: Comparison of traditional In-Context Learning (ICL) where each demonstration example comprises a full text, a question and an answer to DoubleDipper where the demonstrations contain only question-answer pairs, automatically generated on the same input text.
To analyze the performance gain of DoubleDipper compared to the traditional In-Context-Learning, we prompt a representative model from each family (Gemini Pro, Gemma 2B, Llama 7B, and Mistral) with few-shot examples from the same dataset, each example including its own context and question-answer pair.
We report the results in Table 5. For all models, DoubleDipper achieves a higher performance than the traditional ICL. Interestingly, for Gemma 2B and Llama 7B, the traditional ICL not only underperforms DoubleDipper but it also significantly hurts performance compared to the baseline (e.g., -14.3 for Gemma 2B and -14.9 for Llama 7b). This performance drop might be because each demonstration in the traditional ICL setup comprises a few thousand tokens, which exacerbates the models’ existing challenges with processing long-range dependencies.
As mentioned in Section 3, DoubleDipper promotes also efficiency by adding to the original prompt only a few extra tokens, leading to a significantly cheaper inference than the traditional ICL.
How many examples are needed?
In Table 6, we explore the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 3, 5, and 10. Our analysis reveals no significant differences in performance across these variations, while 3 self-generated examples are sufficient to improve performance. This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022b).
$k=3$ $k=5$ $k=10$
Gemini Pro 54.6 54.1 54.7
Gemini Ultra 54.7 54.1 53.8
Gemma-2b-it (v1.1) 41.2 40.8 41.0
Gemma-7b-it (v1.1) 51.3 50.2 49.5
Mistral-7b-instruct (v0.2) 52.8 52.8 52.4
Llama-2-7b-chat 33.3 32.5 31.5
Llama-2-13b-chat 34.9 34.4 33.0
Llama-2-70b-chat 51.6 51.1 51.5
Table 6: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper.
DoubleDipper without identification of supporting paragraphs
To ablate the second principle in DoubleDipper, namely the explicit identification of the supporting paragraphs before generating the answer, we prompt Gemini Pro and Gemini Ultra with self-generated few shot examples that comprise only question-answer pairs (without instructing the model to retrieve the relevant passage(s)). Gemini Pro achieves on average 36.2 points (-18.4) and Gemini Ultra achieves 54.0 (-0.7 points). This indicates that instructing small models (e.g., Gemini Pro) with explicit prediction of supporting paragraphs achieves similar results to large models (Gemini Ultra).
Do few-shot examples really “instruct” the model to comprehend the text or provide clues to the query?
To answer this question, we prompt Gemini Pro and Gemini Ultra with only the self-generated demonstrations without the input context. As expected, removing the input context leads to a huge performance drop, -31.7 for Gemini Pro and -22.8 for Gemini Ultra compared to DoubleDipper. This confirms that DoubleDipper indeed “teaches” the model how to comprehend the input.
Qualitative Analysis
We manually analyze 50 prompts, with few-shot demonstrations produced by Gemini Ultra. Our review confirms that 93.5% of these self-generated QAs are correct, meaning that the question is meaningful and the answer could be found in the corresponding paragraph.
7 Conclusion
We develop DoubleDipper, a straightforward method for enhancing the performance of Question Answering with long context and providing attribution to the relevant paragraph(s) in the input. By recycling the input context to generate the few shot examples, each demonstration includes solely a question, an answer and a pointer to the relevant paragraph, without a separate context, thus effectively addressing the challenging of In-Context-Learning with long context. Experimental results show that our approach ourperforms both the vanilla LLM and the traditional ICL in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
8 Limitations
One notable limitation of our approach is the extended inference time required for generating question-answer pairs. Future research could mitigate this issue by developing smaller, specialized models specifically tailored for QA generation.
Additionally, our evaluation set is constrained to instances that are solely in English and range between 1,000 to 4,000 tokens. Expanding the diversity of languages and token ranges could enhance the robustness and applicability of our findings.
Lastly, although we employ a strategy of randomly sampling $k$ paragraphs from the input to ensure the model engages with varied segments of the text, we did not optimize the selection of these paragraphs. Future work could explore more strategic methods for paragraph selection to potentially enhance the efficacy and relevance of the generated examples.
Appendix A Prompts
Figure 5 shows the zero-shot prompt we use for generating the question-answer pairs in DoubleDipper. For the QA prompts, we use the same instructions and prompt template as the original papers (Lost-in-the-middle and FLenQA) and add a simple line for the instructions in other multi-hop QA datasets: “Please answer the question based on the given passages below.”. For MuSique, since the dataset includes questions that are not answerable, we add the following sentence to the prompt: “If the question can’t be answered given the given passages, please write "unanswerable"”.
basicstyle=,xleftmargin=2.8em,xrightmargin=1.5em, showstringspaces=false, showspaces=false, showtabs=false, tabsize=2, breaklines=true, flexiblecolumns=true, escapeinside=<@@>, breakatwhitespace=true
Figure 5: Template prompt for the generation question-answer pairs.
Appendix B Identification of Supporting Passages
Table 7 presents the F1 results of the tested models for the supporting relevant passages identification. Without any demonstration (zero-shot), all open source models achieve a poor performance, while DoubleDipper significantly improves performance (+37.8 F1 for Llama2 70B, +38 F1 for Gemma 7B, +30.5 F1 for Mistral).
Avg. Lost PIR MonoRel SRT HotPotQA 2Wiki MuSique
Gemini Pro Baseline - - - - - - - -
Zero-shot 54.5 58.2 47.1 75.9 46.6 56.3 58.3 39.4
DoubleDipper (Self) 50.0 59.7 77.4 38.8 54.2 51.2 34.6 34.1
DoubleDipper (PaLM 2) 50.6 60.1 47.3 68.0 21.2 58.8 57.9 41.0
Gemini Ultra Baseline - - - - - - - -
Zero-shot 40.4 57.4 30.3 39.4 17.0 52.4 43.7 42.9
DoubleDipper (Self) 37.5 58.6 25.5 21.5 11.1 55.4 46.4 44.3
DoubleDipper (PaLM 2) 38.4 58.0 28.0 24.6 12.0 55.3 47.0 43.7
Gemma-2b-it (v1.1) Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 23.3 10.4 43.1 37.5 13.5 20.1 13.0 25.3
DoubleDipper (PaLM 2) 25.6 9.1 48.7 48.7 15 18.7 14 25.3
Gemma-7b-it (v1.1) Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 38.0 39.7 44.8 59.8 7.1 41.8 36.5 36.1
DoubleDipper (PaLM 2) 36.9 39.2 45.2 58.6 7.3 40.1 32.8 35
Mistral-7b-instruct (v0.2) Baseline - - - - - - - -
Zero-shot 11.5 6.7 0.7 2.9 0.0 17.2 18.2 34.5
DoubleDipper (Self) 42.0 48.2 66.2 60.6 5.8 41.7 29.3 42.2
DoubleDipper (PaLM 2) 41.5 48.3 70.0 58.6 6.8 40.0 26.0 40.6
Llama-2-7b-chat Baseline - - - - - - - -
Zero-shot 0.8 0.2 0.0 0.0 3.2 0.0 0.0 2.1
DoubleDipper (Self) 19.4 17.5 27.7 20.2 5.7 21.7 19.7 23.3
DoubleDipper (PaLM 2) 19.6 15.8 29.3 21.6 5.7 22.2 20.7 22.0
Llama-2-13b-chat Baseline - - - - - - - -
Zero-shot 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
DoubleDipper (Self) 28.8 16.5 51.0 32.0 18.4 29.6 22.2 31.7
DoubleDipper (PaLM 2) 30.3 13.4 51.6 30.5 15.6 36.4 27.3 36.9
Table 7: Performance (F1) of supporting paragraph(s) prediction. The baseline does not predict supporting paragraphs. DoubleDipper provides a significant performance boost for all open source models, whose performance is close to 0 (except from Mistral) in the zero shot setting.